# Is Mamba Effective for Time Series Forecasting?

### v1

@@ Introduction
We conduct extensive experiments on several real-world datasets, including PEMS (PEMS03, PEMS04, PEMS07, PEMS08), ETT (Electricity Transformer Temperature) (Zhou et al. 2021) (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange-Rate (Lai et al. 2018), Traffic, Weather, Solar-Energy. All of them are multivariate time series datasets adopted by Liu et al. (2023). Tab. 2 shows the statistics of these datasets.
@@
Electricity 321 26,304 1hour
Exchange-Rate 8 7,588 1day
Traffic 862 17,544 1hour
Weather 21 52,696 10min
Solar-Energy 137 52,560 10min
@@
Tab.1 presents a comparative analysis of the overall performance of our models and other baseline models across all datasets. From the data presented in the table, we can deduce the following three observations: (1) The D-Mamba and S-Mamba models generally maintain a performance advantage, albeit a slight one, in most scenarios. (2) D-Mamba outperforms S-Mamba in the majority of cases, suggesting that in TSF tasks, it may be more effective to configure Mamba with varying sensitivities to work collaboratively when extracting relationships between variables. (3) The distinction between S-Mamba and iTransformer lies in their respective approaches to extracting dependencies between variates. On certain datasets, substituting Transformer with Mamba leads to a decrease in effectiveness. This raises an intriguing research question: under what circumstances does Mamba surpass Transformer in performance?
Table 3: Comparison of costs between S/D-Mamba and three baselines, including GPU memory usage, and training time on ECL and Traffic. The number on the left side of the table represents the lookback length $L\in\{96,192,336,720\}$, while the forecasting length $T=96$.
Models D-Mamba S-Mamba iTransformer RLinear Autoformer
Metric Memory Speed Memory Speed Memory Speed Memory Speed Memory Speed
(GiB) (ms/iters) (GiB) (ms/iters) (GiB) (ms/iters) (GiB) (ms/iters) (GiB) (ms/iters)
ECL 96 0.537 39.5 0.402 37.2 0.771 41.5 0.042 18.1 1.696 81.6
192 0.540 40.3 0.411 38.0 0.775 43.6 0.054 22.8 2.426 105.0
336 0.549 41.3 0.427 38.6 0.782 45.5 0.071 30.6 3.702 148.4
720 0.564 43.4 0.467 39.4 0.800 56.8 0.115 47.0 6.991 247.2
Traffic 96 1.667 11.3 1.262 97.3 4.549 161.2 0.085 39.3 1.791 203.2
192 1.676 11.5 1.288 97.5 4.558 154.6 0.115 50.7 2.532 215.2
336 1.725 11.8 1.326 101.5 4.575 160.9 0.161 76.5 3.818 277.9
720 1.749 11.9 1.427 116.6 4.617 167.4 0.279 137.4 7.133 398.3
@@
To evaluate the computational efficiency of the models, We compare the memory usage and computing time of S/D Mamba with several baselines on the ECL and Traffic. Specifically, we conduct independent runs on a single NVIDIA A6000 GPU and meticulously document the results in Tab.3. The numbers in the left column of the table represent the lookback length $L\in\{96,192,336,720\}$ The results indicate that, in comparison to the Transformer-based model, both S-Mamba and D-Mamba exhibit optimization in terms of time cost and GPU memory usage, and demonstrate a performance more akin to that of linear models, R-Linear.
Table 4: Performance comparison between Autoformer, Flashformer and Flowformer and their variants with Mamba encoder block on ECL and Traffic datasets, where we fixed the lookback length $L=96$ and forcasting length $T=96$.
@@
Memory(GiB) 1.59 1.22 0.91 0.63 0.83 0.68 1.78 1.35 1.16 0.80 1.06 0.79
Speed(ms/iters) 76.3 56.1 62.8 40.6 57.4 38.0 126.1 97.3 110.3 91.5 101.5 81.2
@@
Is Mamba a better Variate-Encoder?
In prior studies, the encoder of the Transformer is commonly utilized for variant-encoder tasks, which involve aggregating all variable information at a single time step. From a structural perspective, in most Transformer-based models with encoder-decoder architectures, the self-attention mechanism within the encoder layer can be substituted with the Mamba mechanism. Furthermore, from a utility perspective, Mamba’s selective mechanism can distinguish VC among different variables, enabling it to function as a Variate-Encoder. Therefore, we set an experiment by replacing the encoder layer (self-attention) of Transformer (Vaswani et al. 2017), Autoformer (Wu et al. 2021), Flashformer (Dao et al. 2022) and Flowformer (Wu et al. 2022) with a simple Mamba block to get Trans-M, Auto-M, Fed-M, Flash-M and Flow-M and and compare their performance. The results are astonishing as 4, which straightforward modification with Mamba enhances the performance of these models and also reduces GPU memory usage and running time, which may indicate that Mamba has better results when aggregating multi-variates information.
Can Mamba help learn temporal information?
@@
Prior research has shown that the performance of Transformer-based models does not consistently improve with increasing lookback sequence length $L$, which is somewhat unexpected. A plausible explanation is that the temporal sequence relationship is overlooked under the self-attention mechanism, as it disregards the sequential order, and in some instances, even inverts it. Mamba, resembling a Recurrent Neural Network (Medsker, Jain et al. 2001), concentrates on the preceding window during information extraction, thereby preserving certain sequential attributes. This prompts an exploration of Mamba’s potential effectiveness in temporal sequence information fusion, aiming to address the issue of diminishing or stagnant performance with increasing lookback length. Consequently, we add an additional Mamba block between the encoder Layer and decoder layer of Transformers. The role of this Mamba Block is to add a layer of time sequence dependence from the information output by the encoder layer, to add some information similar to position embedding before the decoder layer processes it. We experiment with Autoformer (Wu et al. 2021), Transformer (Vaswani et al. 2017), and Informer (Zhou et al. 2021) to get Auto-M, Infor-M, and Refor-M, and evaluate their performance with varying lookback lengths. We also test the performance of S/D-Mamba as the lookback length changes. The results are in the Fig.3, from which we can observe three results. (1) S/D-Mamba can enhance its performance as the input lengthens, but we believe this is not solely due to the Mamba Block, but rather to the Variate-based Embedding method proposed by iTransformer (Liu et al. 2023). (2) With the addition of a Mamba Block, all three models show performance improvements. Notably, the performance of Informer has seen the most significant enhancement. (3) Despite these variants’ performance gains sometimes, they do not achieve optimization with longer lookback lengths. This is consistent with the findings of Zeng et al. (2023), which also suggest that encoding temporal sequence information into the model beforehand does not resolve this issue.
@@
Figure 4: Forecasting performance comparison between S/D-Mamba and iTransformer trained on 100% variates with on 40% variates. The lookback length $L=96$ for all datasets. For the PEMS dataset, the output length $T=12$, while for the other datasets $T=96$.
In this vein, iTransformer (Liu et al. 2023) conducts a pivotal experiment. The study involves masking a majority of the variates in a dataset and training the model on a limited subset of variates. Subsequently, the model was tasked with forecasting all variates, including those previously unseen, based on the learned information from the few variates. Building on this, we seek to evaluate the generalization capabilities of Mamba in TSF tasks. An experiment is proposed wherein the S/D-Mamba are trained on merely 40% of the variates in the PEMS03, PEMS04, ECL, Weather, Traffic, and Solar datasets. Then they are employed to predict 100% variates, and the results are subjected to statistical analysis. The outcomes of this investigation in 4reveal that both S-Mamba and D-Mamba exhibit generalization potential in the six datasets, which proves their generalizability in TSF tasks. However, under this training paradigm, they demonstrate inferior performance compared to iTransformer, underscoring their generalizability in TSF tasks.

### v2

@@ Introduction
We conduct extensive experiments on several real-world datasets, including Electricity, Exchange-Rate (Lai et al. 2018), Traffic, Weather, Solar-Energy, PEMS (PEMS03, PEMS04, PEMS07, PEMS08), ETT (Electricity Transformer Temperature) (Zhou et al. 2021) (ETTh1, ETTh2, ETTm1, ETTm2). All of them are multivariate time series datasets adopted by Liu et al. (2023). Tab. 1 shows the statistics of these datasets.
@@
Electricity 321 26,304 1hour
Exchange-Rate 8 7,588 1day
Traffic 862 17,544 1hour
Weather 21 52,696 10min
Solar-Energy 137 52,560 10min
@@
Table 2: Full results of S-Mamba and balines. The lookback length $L$ is set to 96 and the forecast length $T$ is set to 12, 24, 48, 96 for PEMS and 96, 192, 336, 720 for other datasets. Results are averaged from all forecast lengths and full Results are in the Appendix A. The best results are highlighted in bold red font, while the second best results are presented in underlined blue font. The results of baselines are reported by Liu et al. (2023).
Models S-Mamba iTransformer RLinear PatchTST Crossformer TiDE TimesNet DLinear FEDformer Autoformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL 0.170 0.265 0.178 0.270 0.219 0.298 0.205 0.290 0.244 0.334 0.251 0.344 0.192 0.295 0.212 0.300 0.214 0.327 0.227 0.338
Exchange 0.367 0.408 0.360 0.403 0.378 0.417 0.367 0.404 0.940 0.707 0.370 0.413 0.416 0.443 0.354 0.414 0.519 0.429 0.613 0.539
Traffic 0.414 0.276 0.428 0.282 0.626 0.378 0.481 0.304 0.550 0.304 0.760 0.473 0.620 0.336 0.625 0.383 0.610 0.376 0.628 0.379
Weather 0.251 0.276 0.258 0.278 0.272 0.291 0.259 0.281 0.259 0.315 0.271 0.320 0.259 0.287 0.265 0.317 0.309 0.360 0.338 0.382
Solar-Energy 0.240 0.273 0.233 0.262 0.369 0.356 0.270 0.307 0.641 0.639 0.347 0.417 0.301 0.319 0.330 0.401 0.291 0.381 0.885 0.711
PEMS03 0.122 0.228 0.113 0.221 0.495 0.472 0.180 0.291 0.169 0.281 0.326 0.419 0.147 0.248 0.278 0.375 0.213 0.327 0.667 0.601
PEMS04 0.103 0.211 0.111 0.221 0.526 0.491 0.195 0.307 0.209 0.314 0.353 0.437 0.129 0.241 0.295 0.388 0.231 0.337 0.610 0.590
PEMS07 0.089 0.188 0.101 0.204 0.504 0.478 0.211 0.303 0.235 0.315 0.380 0.440 0.124 0.225 0.329 0.395 0.165 0.283 0.367 0.451
PEMS08 0.148 0.224 0.150 0.226 0.529 0.487 0.280 0.321 0.268 0.307 0.441 0.464 0.193 0.271 0.379 0.416 0.286 0.358 0.814 0.659
ETTm1 0.398 0.405 0.407 0.410 0.414 0.407 0.387 0.400 0.513 0.496 0.419 0.419 0.400 0.406 0.403 0.407 0.448 0.452 0.588 0.517
ETTm2 0.288 0.332 0.288 0.332 0.286 0.327 0.281 0.326 0.757 0.610 0.358 0.404 0.291 0.333 0.350 0.401 0.305 0.349 0.327 0.371
ETTh1 0.455 0.450 0.454 0.447 0.446 0.434 0.469 0.454 0.529 0.522 0.541 0.507 0.458 0.450 0.456 0.452 0.440 0.460 0.496 0.487
ETTh2 0.381 0.405 0.383 0.407 0.374 0.398 0.387 0.407 0.942 0.684 0.611 0.550 0.414 0.427 0.559 0.515 0.437 0.449 0.450 0.459
@@
Tab.2 presents a comparative analysis of the overall performance of our models and other baseline models across all datasets. From the data presented in the table, we summarize two observations and attach the analysis: (1) S-Mamba has attained commendable outcomes on the Electricity, Traffic, PEMS, PEMS and Solar datasets. These datasets are distinguished by their numerous variates, most of which have the same period. It is worth noting that variates with the same period are more likely to contain learnable VC. Mamba VC Fusion Layer benefits from this characteristic and improves S-Mamba performance. (2) In the context of the ETT, and Exchange data sets, S-Mamba did not demonstrate a pronounced superiority in performance; indeed, it exhibited a suboptimal outcome. This can be attributed to the fact that these datasets are characterized by a limited number of variates, predominantly of an aperiodic nature. Consequently, there are weak VCs between these variates, and the employment of VC by S-Mamba may inadvertently introduce noise into the predictive model, thus impeding its predictive accuracy. (3) The Weather data set is special in that it has fewer variables and most variables are aperiodic, but S-Mamba still achieves the best performance on it.
Refer to captionRefer to captionRefer to captionRefer to caption
(a) S-Mamba
Refer to captionRefer to captionRefer to captionRefer to caption
(b) iTransformer
Figure 3: Comparison of forecasts between S-Mamba and iTransformer on four datasets when the input length is 96 and the forecast length is 96. The blue line represents the ground truth and the red line represents the forecast.
Furthermore, to provide a more intuitive assessment of S-Mamba’s forecast capabilities, we visually compare the predictions of S-Mamba and the leading baseline, iTransformer, on four datasets: Electricity, Weather, Traffic, and Exchange, through graphical representation. Specifically, we randomly select a variate and then input its lookback sequence, where the true subsequent sequence is depicted as a blue line and the model’s forecast is represented by a red line in Fig.3. It is evident that on the ECL, Weather, and Traffic datasets, S-Mamba’s predictions closely approximate the actual values, with nearly perfect alignment observed on the ECL and Traffic datasets. However, on the Exchange dataset, both models exhibit relatively modest performance.
(a) Electricity
(b) Traffic
Figure 4: Comparison of S-Mamba and five baselines on MSE, Training Time, and GPU Memory.
@@
To evaluate the computational efficiency of the models, We compare the memory usage and computing time of S-Mamba with several baselines on the ECL and Traffic. Independent runs are conducted on a single NVIDIA RTX3090 GPU and meticulously document the results in Fig.4. In our analysis, a bubble chart to depict the measurement outcomes, wherein the vertical axis denotes the Mean Squared Error (MSE), the horizontal axis quantifies the training duration, and the bubble magnitude correlates with the video memory utilization. The visualization reveals that the S-Mamba algorithm attains the most favorable MSE metric across both the Electricity and Traffic datasets. When benchmarked against Transformer-based models, S-Mamba typically necessitates a reduced training time while concurrently minimizing GPU memory consumption. While the RLinear model does utilize minimal GPU memory and curtails training time, it does not confer a competitive edge in terms of predictive precision. Overall, S-Mamba manifests exemplary predictive accuracy with a low computational resource footprint.
Table 3: Ablation study on ECL, Traffic, Weather, and Solar-Energy. The lookback length $S=96$, while the forecast length $T\in\left\{96,192,336,720\right\}$. Results are averaged from all forecast lengths.
Design Variate Temporal ECL Traffic Weather Solar-Energy
MSE MAE MSE MAE MSE MAE MSE MAE
S-Mamba Mamba FFN 0.170 0.265 0.411 0.273 0.251 0.276 0.240 0.273
Replace Mamba uni-Mamba 0.182 0.280 0.538 0.342 0.244 0.268 0.250 0.282
Mamba bi-Mamba 0.182 0.281 0.551 0.371 0.246 0.271 0.255 0.282
Mamba Attention 0.179 0.278 0.561 0.379 0.247 0.272 0.261 0.288
Attention FFN 0.178 0.270 0.428 0.282 0.258 0.279 0.233 0.262
w/o Mamba w/o 0.174 0.269 0.418 0.278 0.253 0.279 0.247 0.278
w/o FFN 0.193 0.276 0.461 0.294 0.265 0.283 0.261 0.283
Ablation Study
To evaluate the efficacy of the components within S-Mamba, we conduct ablation studies by substituting or eliminating the VC and TD encoding layers. Specifically, the VC encoding layer was replaced with an Attention mechanism or entirely removed. This modification is predicated on the empirical evidence from iTransformer experiments (Liu et al. 2023), which demonstrate that Attention was the optimal encoder for VC. Conversely, the TD encoding layer is replaced with Attention, bidirectional Mamba, unidirectional Mamba, or omitted altogether. The choice of bidirectional Mamba, which is set to benchmark Attention, is made to facilitate global temporal information extraction. The rationale behind employing unidirectional Mamba is its resemblance to RNN models, so inherently possesses the capacity to preserve sequential relationships, thereby making it a suitable candidate for evaluating the impact of sequential encoding on TD.
Our experimental investigations are conducted on four datasets: Electricity, Traffic, Weather, and Solar Energy. The findings from these experiments in Tab.6 indicate that Mamba exhibite superior performance in VC encoding, whereas the Feed-Forward Network (FFN) maintained its dominance in TD encoding.
Can Variate Order Affect Mamba’s Performance?
In the context of the S-Mamba model, the existence of the Hippo matrix (Gu et al. 2020) theoretically enables the retention of information over an indefinite length. However, this very feature predisposes the model to initially prioritize proximal information. S-Mamba employs a bidirectional Mamba layer to extract VC. This initial bias towards neighboring variates could potentially impede the acquisition of a global VC, especially when relevant variates are distant from each other. This consideration prompted us to investigate the impact of variate ordering on the performance of S-Mamba.
We hypothesize that within the dataset, variates are more likely to extract valuable information from other periodic variates. Conversely, the propensity to acquire noise increases when interacting with aperiodic variates. To address this, we apply the Fourier transform (Bracewell 1989) to help categorize variates into periodic and aperiodic groups. Our experimentation focuses on the ECL dataset, notable for its numerous periodic and aperiodic variates. The initial arrangement of these variates is depicted in the accompanying Fig.5.
Subsequent trials involve repositioning the aperiodic variates towards the middle or end of the variates sequence, followed by evaluating the predictive capabilities of the models trained on these modified datasets. For comparative analysis, we also included experiments with the iTransformer. The variate distribution and corresponding model performance are illustrated in the Fig. . Our findings suggest that the S-Mamba model’s performance remains largely unaffected by the perturbation of variate order. This implies that through adequate training, the S-Mamba can effectively mitigate the initial limitation of the Hippo matrix, which biases attention towards nearby information.
Table 4: Performance comparison between Autoformer, Flashformer and Flowformer and their variants with Mamba encoder block on ECL and Traffic datasets, where we fixed the lookback length $L=96$ and forecasting length $T=96$.
@@
Memory(GiB) 1.59 1.22 0.91 0.63 0.83 0.68 1.68 1.35 1.16 0.80 1.06 0.79
Speed(ms/iters) 54.6 45.1 62.8 40.6 57.4 38.0 126.1 97.3 110.3 91.5 101.5 81.2
@@
Can Mamba Outperform Advanced Transformers?
Beyond the foundational Transformer architecture, some advanced Transformers have been introduced, predominantly focusing on the augmentation of the self-attention mechanism. We aim to determine whether Mamba can still maintain an advantage over these advanced Transformers. To this end, we conduct a comparative experiment in which we directly replace the Encoder layer of three advanced transformers: Autoformer (Wu et al. 2021), Flashformer (Dao et al. 2022) and Flowformer (Wu et al. 2022) with a unidirectional Mamba for TSF tasks to get Trans-M, Auto-M, Fed-M, Flash-M and Flow-M and and compare their performance. The results are astonishing as Tab.4, which straightforward modification with Mamba enhances the performance of these models and also reduces GPU memory usage and running time.
Figure 5: In the ECL dataset, adjust the distribution of periodic and aperiodic variates. The left side represents the distribution, while the right side indicates models’ performance.
Can S-Mamba Benefit from Increasing Lookback Length?
@@
Prior research has shown that the performance of Transformer-based models does not consistently improve with increasing lookback sequence length $L$, which is somewhat unexpected. A plausible explanation is that the temporal sequence relationship is overlooked under the self-attention mechanism, as it disregards the sequential order, and in some instances, even inverts it. Mamba, resembling a Recurrent Neural Network (Medsker, Jain et al. 2001), concentrates on the preceding window during information extraction, thereby preserving certain sequential attributes. This prompts an exploration of Mamba’s potential effectiveness in temporal sequence information fusion, aiming to address the issue of diminishing or stagnant performance with increasing lookback length. Consequently, we add an additional Mamba block between the encoder Layer and decoder layer of Transformers. The role of this Mamba Block is to add a layer of time sequence dependence from the information output by the encoder layer, to add some information similar to position embedding before the decoder layer processes it. We experiment with Autoformer (Wu et al. 2021), Transformer (Vaswani et al. 2017), and Informer (Zhou et al. 2021) to get Auto-M, Infor-M, and Refor-M, and evaluate their performance with varying lookback lengths. We also test the performance of S-Mamba as the lookback length changes. The results are in the Fig.6, from which we can observe three results. (1) S-Mamba can enhance its performance as the input lengthens, but we believe this is not solely due to the Mamba Block, but rather to the Variate-based Embedding method proposed by iTransformer (Liu et al. 2023). (2) With the addition of a Mamba Block, all three models show performance improvements. Notably, the performance of Informer has seen the most significant enhancement. (3) Despite these variants’ performance gains sometimes, they do not achieve optimization with longer lookback lengths. This is consistent with the findings of Zeng et al. (2023), which also suggest that encoding temporal sequence information into the model beforehand does not resolve this issue.
@@
Figure 7: Forecasting performance comparison between S-Mamba and iTransformer trained on 100% variates with on 40% variates. The lookback length $L=96$ for all datasets. For the PEMS dataset, the output length $T=12$, while for the other datasets $T=96$.
In this vein, iTransformer (Liu et al. 2023) conducts a pivotal experiment. The study involves masking a majority of the variates in a dataset and training the model on a limited subset of variates. Subsequently, the model was tasked with forecasting all variates, including those previously unseen, based on the learned information from the few variates. Building on this, we seek to evaluate the generalization capabilities of Mamba in TSF tasks. An experiment is proposed wherein the S-Mamba are trained on merely 40% of the variates in the PEMS03, PEMS04, ECL, Weather, Traffic, and Solar datasets. Then they are employed to predict 100% variates, and the results are subjected to statistical analysis. The outcomes of this investigation in 7 reveal that S-Mamba exhibits generalization potential in the six datasets, which proves their generalizability in TSF tasks.